OpenStudy (anonymous):

A right triangle has a hypotenuse of 70 feet and a leg of 35 feet. What is the length of the other leg?

Parth (parthkohli):

The Pythagorean \(35^2 + b^2 = 70^2\)

OpenStudy (anonymous):

\[\large a^2 + b^2 = c^2\] here c is the hypotenuse. \[a^2 + 35^2 = 70^2\] Find a here..

Parth (parthkohli):

Gotta solve for b right there :)

OpenStudy (jiteshmeghwal9):

use pythagoras theorem\[(h)^2=(p)^2+(s)^2\]

OpenStudy (anonymous):

use pythagorean theorem... a^2 + b^2 = c^2 where c is the hypotenuse.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

A. 51 feet B. 61 feet C. 78 feet D. 80 feet

OpenStudy (anonymous):

try to solve it. It will help you a lot.

OpenStudy (anonymous):

Find a there you will get one of these values..

OpenStudy (anonymous):

ok i am

OpenStudy (anonymous):

ok got it thanks